Soybean meal is 12% protein; Cornmeal is 6% protein. How many pounds of each should be mixed together to get a 240 pound mixture that is 7% protein?
Answer by checkley77(12569)

.12x+.06(240-x)=.07*240
.12x+14.4-.06x=16.8
.06x=16.8-14,4
.06x=2.4
x=2.4/.06
x=40 pounds of soybeans was used.
240-40=200 pounds of cornmeal was used.
Proof:
.12*40+.06*200=16.8
4.8+12=16.8
16.8=16.8